OpenStudy (anonymous):

question: if 6delta x*x+3deltax^2/ delta x then can we factor and cancel delta x?

OpenStudy (anonymous):

\[\frac{ 6deltax*x +3deltax^2 }{ deltax }\]

OpenStudy (anonymous):

example: can we factor delta x out? and be left with 6x+3deltax?

OpenStudy (anonymous):

or does the delta x next to the 3 also get cancelled

OpenStudy (anonymous):

@Psymon @genius12 @dumbcow @satellite73

OpenStudy (dumbcow):

no you are correct...i am assuming its (delta x)^2 right?

OpenStudy (anonymous):

nah it's exactly how i wrote it. i was trying to simply and i was wondering if there a delta x inside the parenthesis gets cancelled by crossing it out or does deltax=0 and so we only work with 6 only (x) which (x=-6) which = -36

OpenStudy (dumbcow):

does this have to do with limits and definition of derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathcalculus Yes you can.

OpenStudy (anonymous):

\[\bf \frac{ 6x \Delta x+3(\Delta x)^2 }{ \Delta x }=\frac{ 3 \cancel{\Delta x}(2x+ \Delta x) }{ \cancel{\Delta x} }=3(2x + \Delta x)\]

OpenStudy (anonymous):

@mathcalculus

OpenStudy (anonymous):

thank you @genius12